OpenStudy (anonymous):

WILL FAN AND MEDAL FOR BEST ANSWER: 5x-y=13, 2x+y=15

OpenStudy (malcolmmcswain):

I assume you are looking for the solution to this system of equations?

OpenStudy (anonymous):

Yes.

OpenStudy (malcolmmcswain):

The answer is (4,7)

OpenStudy (malcolmmcswain):

Allow me to explain.

OpenStudy (anonymous):

5(3) - (2)=13

OpenStudy (malcolmmcswain):

5*4-7=13

OpenStudy (anonymous):

Could you maybe show me?

OpenStudy (malcolmmcswain):

5*4-7=13 2*4+7=15

OpenStudy (malcolmmcswain):

Do you see now?

OpenStudy (anonymous):

2(5)+5=15

OpenStudy (anonymous):

Add the two equations together, the y terms will subtract out. 7x = 28 x = 4

OpenStudy (anonymous):

Okay keep going.

OpenStudy (malcolmmcswain):

That's all you need.

OpenStudy (malcolmmcswain):

Using x, you can sole for y.

OpenStudy (anonymous):

@malcolmmcswain that isn't helpful like at all

OpenStudy (malcolmmcswain):

5*4-y=13, solve for y.

OpenStudy (malcolmmcswain):

I'm sorry. Do you know how to solve an algebraic equation?

OpenStudy (anonymous):

Then 8 + y = 15 y = 7

OpenStudy (malcolmmcswain):

I can go through step by step.

OpenStudy (anonymous):

No that's why I need your help.

OpenStudy (malcolmmcswain):

Looks like @robtobey has got it.

OpenStudy (malcolmmcswain):

Do you understand now?

OpenStudy (anonymous):

Yes. Thanks for your effort though. I'll give you a medal for trying.

OpenStudy (malcolmmcswain):

Thanks! I appreciate it. I'm glad you understand now.

OpenStudy (anonymous):

No problem.